# Learning to Optimize for Mixed-Integer Non-linear Programming

### v3

@@ 1 Introduction
Mixed-integer optimization is fundamental to a broad spectrum of real-world applications spanning problems in fields as diverse as pricing Kleinert et al. (2021), battery dispatch (Nazir & Almassalkhi, 2021), transportation (Schouwenaars et al., 2001), and optimal control (Marcucci & Tedrake, 2021). These problems involve discrete decisions, such as determining the number of items or the activation of generators, combined with complex non-linear system constraints. Mixed-integer linear programming (MILP) has been widely adopted due to its well-established solution techniques. However, many practical problems exhibit non-linear relationships, leading to mixed-integer non-linear programs (MINLPs). Unlike MILPs, where techniques such as branch-and-bound (Land & Doig, 2010), cutting planes (Gomory, 2010), and heuristics (Crama et al., 2005; Johnson & McGeoch, 1997) have matured, MINLPs require more complex approaches due to the combination of discrete variables and non-convex constraints and objective function. Standard methods include outer approximation (Fletcher & Leyffer, 1994), spatial branch-and-bound (Belotti et al., 2009), and decomposition techniques (Nowak, 2005), but these often struggle to scale to large problems.
Many applications demand that MINLPs be solved within a limited time budget, further complicating the picture. To overcome this, learning-to-optimize (L2O) methods offer a promising alternative by leveraging machine learning (ML) to enhance or even replace conventional optimization approaches. In particular, end-to-end optimization directly maps input instance parameters to solutions of optimization problems through a trained model (Kotary et al., 2021a; Chen et al., 2022). By identifying patterns in a distribution of similar instances of the same optimization problem and predicting solutions accordingly, end-to-end optimization can bypass traditional, computationally intensive optimization methods, enabling faster computation and improved scalability.
@@
This work tackles, for the first time, the non-differentiability associated with predicting integer variables using a deep neural network, in conjunction with non-linear objective and constraints. This challenge has been underexplored in learning-based methods due to the absence of useful gradient information. To that end, we propose two differentiable correction layers for rounding, allowing for gradient-based optimization of a neural network that generates high-quality integer solutions while maintaining feasibility. Our contributions are as follows:
  • We study the learning-to-optimize problem in the context of parametric MINLP, enabling, for the first time, quick, dynamic solution generation as problem parameters change.
  • We propose two novel differentiable correction layers that effectively handle the non-differentiability of integer outputs in deep learning models. Combined with a penalty method for soft constraint satisfaction, we are able to learn a neural network mapping from instance parameters to solutions through gradient-based learning. Not only does this yield an extremely fast heuristic at test time, but the method is also self-supervised and thus efficiently trainable. Additionally, we introduce a strategy of progressively increasing penalties to further improve feasibility in the generated solutions.
  • We conduct extensive experiments on three problem classes from the literature: a convex integer quadratic problem, a non-convex integer problem, and a non-convex mixed-integer problem. Our learning-based methods consistently yield high-quality, feasible solutions extremely fast, outperforming exact solvers and heuristics. For the largest test instances we consider, the baselines fail to produce any solutions, whereas our methods continue to generate good, feasible solutions in most instances.
@@ 2 Related Work
End-to-end optimization focuses on training machine learning models to predict the problem solutions, bypassing the need for computationally expensive solvers. One of the early approaches was proposed by Hopfield & Tank (1985), who used Hopfield networks to solve the traveling salesperson problem by incorporating a Lagrangian penalty for constraint feasibility. Similarly, Fioretto et al. (2020) applied the Lagrangian penalty in the context of continuous non-linear optimization for energy systems. In addition to penalty-based methods for ensuring feasibility, Pan et al. (2020) embedded certain constraints directly into neural networks by leveraging the range of output values and solving linear systems. Although these supervised learning methods significantly reduce inference time, they typically require large offline datasets of solutions (Gleixner et al., 2021; Kotary et al., 2021b), which can be impractical for large-scale problems where generating solutions is computationally expensive. This limitation highlights the need for self-supervised learning approaches (Donti et al., 2021), which minimize both the objective function and constraint violation from the predicted values, without relying on the imimitation of pre-solved solutions. Our method first extends this self-supervised paradigm to problems involving discrete decision variables, further broadening its applicability to mixed-integer optimization.
@@
There has been significant interest in using ML to accelerate the solution of integer programs. The vast majority of the work in this space focuses on learning search strategies for exact MILP solvers. This includes parameter tuning (Xu et al., 2011), preprocessing (Berthold & Hendel, 2021), branching variable selection (Khalil et al., 2016; Alvarez et al., 2017; Gasse et al., 2019; Zarpellon et al., 2021), node selection (He et al., 2014), heuristic selection (Chmiela et al., 2021), and cut selection and generation (Deza & Khalil, 2023). Another line of research in ML-for-MILP relates to learning to generate integer solutions heuristically (Nair et al., 2020; Khalil et al., 2022; Ding et al., 2020; Sonnerat et al., 2021; Song et al., 2020; Bertsimas & Stellato, 2022; Huang et al., 2023; Ye et al., 2024). We refer to the surveys of Bengio et al. (2021) and Zhang et al. (2023) for more details. In contrast, there has been much less work on MINLP. Illustrative examples include the work of Cauligi et al. (2022) who proposed a two-stage algorithm for quickly finding high-quality solutions for mixed-integer convex programs (MICPs), Baltean-Lugojan et al. (2019) who use supervised learning to select cuts for quadratic optimization, Nowak et al. (2018) who learn to solve quadratic assignment problems with graph networks, and Bonami et al. (2022) who use a classifier to decide on the linearization of mixed-integer quadratic problems. Most relevant to our method is the recently proposed SurCO approach of Ferber et al. (2023). They focus on mixed-integer problems with non-linear objective and linear constraints, learning to approximate the former with a linear function for a simpler heuristic optimization. Our approach differs from all of the above in its scope, addressing the most general class of MINLPs.
@@ 3 Learning to Optimize MINLPs: a Problem Formulation
$\min_{\Theta}\quad$ $\mathbb{E}_{{\bm{\xi}}\sim\mathcal{P}_{\xi}}\,{\bf f}({\bf x^{\bm{\xi}}},{\bm{\xi}})$ (1a)
s.t. ${\bf g}({\bf x^{\bm{\xi}}},{\bm{\xi}})\leq 0,$ $\forall{{\bm{\xi}}\in\mathcal{P}_{\xi}}$ (1b)
${\bf x^{\bm{\xi}}}\in\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}},$ $\forall{{\bm{\xi}}\in\mathcal{P}_{\xi}}$ (1c)
${\bf x^{\bm{\xi}}}=\bm{\psi}_{\Theta}({\bm{\xi}}),$ $\forall{{\bm{\xi}}\in\mathcal{P}_{\xi}}$ (1d)
Here, $\mathcal{P}_{\xi}$ represents the distribution over parametric MINLP instances and ${\bm{\xi}}\in\mathbb{R}^{n_{\xi}}$ is a vector of instance parameters. Vector ${\bf x^{\bm{\xi}}}\in\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}}$ represents the mixed-integer decision variables for parameters ${\bm{\xi}}$. The mapping $\bm{\psi}_{\Theta}({\bm{\xi}})$ is a neural network with weights $\Theta$ that outputs a parametric solution ${\bf x^{\bm{\xi}}}$ for parameters ${\bm{\xi}}$. The goal is to find the neural network weights that minimize the expected objective function value in eq. 1a over the parameter distribution, subject to the constraints in eq. 1b. Note that $\bf{g}(\cdot)$ is a vector-valued function representing one or more inequality constraints. As is typical in MINLP, we assume that the objective and constraint functions are differentiable.
As is typical, we will train the neural network using empirical risk minimization on a sample of $m$ training instances drawn from the parameter distribution $\mathcal{P}_{\xi}$. Then, the average value of the objective function ${\bf f}(\cdot)$ serves as a natural loss function. Our approach is self-supervised since the loss calculation does not require any labeled data. This is particularly appealing as computing optimal or even feasible solutions to a MINLP is in general extremely challenging. Solely minimizing the average objective is insufficient if the solutions violate the constraints. Therefore, similarly to Donti et al. (2021), we incorporate penalty terms into the loss function to account for constraint violations, enhancing the feasibility of the solution and resulting in a soft-constrained empirical risk minimization loss function given:
$\mathcal{L}(\Theta)=\frac{1}{m}\sum_{i=1}^{m}\Big{(}{\bf f}({\bf x}^{i},{\bm{\xi}}^{i})+\lambda\cdot\|\left({\bf g}({\bf x}^{i},{\bm{\xi}}^{i})\right)_{+}\|_{2}^{2}\Big{)}\text{ with }{\bf x}^{i}=\bm{\psi}_{\Theta}({\bm{\xi}}^{i})\;\quad\forall i\in[m],$ (2)
where $m$ is the number of sampled parametric scenarios in the training dataset, $\left(\cdot\right)_{+}$ ensures only positive constraint violations are penalized (implemented via a ReLU function), $\|\cdot\|_{2}$ denotes the Euclidean norm, and $\lambda>0$ is a hyperparameter that balances the trade-off between minimizing the objective function and satisfying the constraints.
@@ 5 Learning to Optimize MINLPs with Correction Layers
Figure 1: Conceptual diagram for our method. self-supervised differentiable programming-based solution approach for parametric MINLP problems.
@@
1:Training instance ${{\bm{\xi}}^{i}}$, neural networks $\pi_{\Theta_{1}}(\cdot)$ and $\delta_{\Theta_{2}}(\cdot)$
@@
We are now ready to introduce the two differentiable correction layers designed to handle the integrality constraints of MINLPs. They differ in how they round an integer variable’s fractional value but are equally easy to train with gradient descent and fast at test time. We decompose the mapping $\psi_{\Theta}:\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}}$ from an instance parameter vector to a candidate mixed-integer solution into two steps.
  1. 1.
  2. 2.
    The second step is a differentiable correction layer $\varphi_{\Theta_{2}}:\mathbb{R}^{n_{r}+n_{z}}\times\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}}$ that takes as input the instance parameter vector and the continuous solution produced in the first step, and outputs a candidate mixed-integer solution while maintaining differentiability. Here, $\Theta_{2}$ represents the weights of the neural network $\delta_{\Theta_{2}}:\mathbb{R}^{n_{r}+n_{z}}\times\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}+n_{z}}$, which is applied within the correction layer $\varphi_{\Theta_{2}}$.
Finally, during training, the loss function eq. 2 is applied to the neural network weights $\Theta=\Theta_{1}\cup\Theta_{2}$, implicitly taking into account the objective function value and constraint violations of the predicted mixed-integer solution $\hat{\bf x}^{i}$. This process is illustrated in Figure 1.
The following subsections describe two distinct approaches for designing the correction layer $\varphi_{\Theta_{2}}$; the same network $\pi_{\Theta_{1}}$ is used in both approaches. Algorithm 1 summarizes both of our approaches. Line 1 invokes the first step’s network $\pi_{\Theta_{1}}$ and lines 2–11 describe both versions of $\varphi_{\Theta_{2}}$. Our correction layers are simple and efficiently computable. While the STE and Gumbel-Sigmoid techniques have been used to train binarized or quantized neural networks, they have not been leveraged in the context of learning-to-optimize to our knowledge. As we will see in the experimental results, the simplicity of the correction layers is key to fast solution generation in large-scale MINLP problems.
@@
Rounding Classification.
Line 6 of Algorithm 1 is the key step in the rounding classification (RC) approach. For the integer variables of a problem instance, denoted with the subscript $z$, RC applies a stochastic soft-rounding to the output ${\bf h}^{i}_{z}$ of the neural network $\delta_{\Theta_{2}}(\bar{\bf x}^{i},{\bm{\xi}}^{i})$, yielding ${\bf b}^{i}\in\{0,1\}^{n_{z}}$. An entry of ${\bf b}^{i}$ determines whether the continuously relaxed value $\bar{\bf x}^{i}_{z}$ of the corresponding variable is rounded down or up. In the backward pass, STE is used in line 4 for the rounding down operation. In line 6, the derivative of the Sigmoid function is used.
Learnable Threshold.
The key steps of the learnable threshold (LT) approach are described in lines 8 and 9 of Algorithm 1. Rather than use Gumbel-Sigmoid for the rounding as in RC, LT learns a vector of per-variable rounding thresholds, ${\bf v}^{i}\in[0,1]^{n_{z}}$, that is generated by the Sigmoid in line 8 of Algorithm 1. A variable is rounded up if the fractional part of its relaxed value, $(\bar{{\bf x}}^{i}_{z}-\hat{{\bf x}}^{i}_{z})$, exceeds the threshold. The indicator function $\mathbb{I}(\cdot)$ in line 9 produces a binary output in the forward pass. In the backward pass, the gradient is approximated by that of the Sigmoid function with a slope:
Here, the slope is set to 10 to sharpen the Sigmoid function.
Illustration.
Figure 2: The relaxed solutions $\bar{x},\bar{y}$ and the rounding solutions $\hat{x},\hat{y}$ across different epochs of training for the same sample instance using the Rounding Classification approach.
$\min_{x\in\mathbb{R},y\in\mathbb{Z}}$ $\quad(a-x)^{2}+50(y-x^{2})^{2}$
In this formulation, $x$ is a continuous decision variable and $y$ is an integer decision variable, subject to linear constraints. The instances have parameters $a$ and $b$, which represent the input features to the neural network; for the instance illustrated in Figure 2, these are set to $3.83$ and $6.04$, respectively.
Ablation Study.
The two baseline methods isolate parts of our full pipeline, serving as ablations to clarify the role of our correction layers. In Rounding after Learning (RL), rounding is applied to the relaxed solution generated by the learning-based methods after training is complete, meaning the rounding step does not participate in training. In Rounding with STE (RS), the continuous values predicted by the first neural network $\pi_{\Theta_{1}}$ are rounded to nearest integer during training using the Straight-Through Estimator (STE), enabling gradients to pass through the rounding operator, though this rounding is not learnable. For further details, refer to Appendix C.

### v4

@@ 1 Introduction
Mixed-integer optimization is fundamental to a broad spectrum of real-world applications spanning problems in fields as diverse as pricing Kleinert et al. (2021), battery dispatch (Nazir & Almassalkhi, 2021), transportation (Schouwenaars et al., 2001), and optimal control (Marcucci & Tedrake, 2020). These problems involve discrete decisions, such as determining the number of items or the activation of generators, combined with complex non-linear system constraints. Mixed-integer linear programming (MILP) has been widely adopted due to its well-established solution techniques. However, many practical problems exhibit non-linear relationships, leading to mixed-integer non-linear programs (MINLPs). Unlike MILPs, where techniques such as branch-and-bound (Land & Doig, 2010), cutting planes (Gomory, 2010), and heuristics (Crama et al., 2005; Johnson & McGeoch, 1997) have matured, MINLPs require more complex approaches due to the combination of discrete variables and non-convex constraints and objective function. Standard methods include outer approximation (Fletcher & Leyffer, 1994), spatial branch-and-bound (Belotti et al., 2009), and decomposition techniques (Nowak, 2005), but these often struggle to scale to large problems.
Many applications demand that MINLPs be solved within a limited time budget, further complicating the picture. To overcome this, learning-to-optimize (L2O) methods offer a promising alternative by leveraging machine learning (ML) to enhance or even replace conventional optimization approaches. In particular, end-to-end optimization directly maps input instance parameters to solutions of optimization problems through a trained model (Kotary et al., 2021b; Chen et al., 2022a). By identifying patterns in a distribution of similar instances of the same optimization problem and predicting solutions accordingly, end-to-end optimization can bypass traditional, computationally intensive optimization methods, enabling faster computation and improved scalability.
@@
This work tackles, for the first time, the non-differentiability associated with predicting integer variables using a deep neural network, in conjunction with non-linear objective function and constraints. This challenge has been underexplored in learning-based methods due to the absence of useful gradient information. To that end, we propose two differentiable correction layers for rounding, allowing for gradient-based optimization of a neural network that generates high-quality integer solutions while maintaining feasibility. Our contributions are as follows:
  • We initiate the study of the learning-to-optimize problem in MINLP for the first time in the literature, a paradigm that can enable efficient solution generation as problem parameters vary.
  • We develop differentiable correction layers that perform soft rounding of neural network outputs into integer assignments to decision variables.
  • We adopt a self-supervised approach that requires no labeled data for training, making our method efficient and scalable to large problem instances.
  • We evaluate our methods on diverse problem benchmarks and show that they find high-quality solutions extremely fast even for large-scale instances where other methods fail.
@@ 2 Related Work
End-to-end optimization focuses on training machine learning models to predict the problem solutions, bypassing the need for computationally expensive solvers. One of the early approaches was proposed by Hopfield & Tank (1985), who used Hopfield networks to solve the traveling salesperson problem by incorporating a Lagrangian penalty for constraint feasibility. Similarly, Fioretto et al. (2020) applied the Lagrangian penalty in the context of continuous non-linear optimization for energy systems. In addition to penalty-based methods for ensuring feasibility, Pan et al. (2020) embedded certain constraints directly into neural networks by leveraging the range of output values and solving linear systems. Although these supervised learning methods significantly reduce inference time, they typically require large offline datasets of solutions (Gleixner et al., 2021; Kotary et al., 2021a), which can be impractical for large-scale problems where generating solutions is computationally expensive. This limitation highlights the need for self-supervised learning approaches (Donti et al., 2021), which minimize both the objective function and constraint violation from the predicted values, without relying on the imimitation of pre-solved solutions. Our method first extends this self-supervised paradigm to problems involving discrete decision variables, further broadening its applicability to mixed-integer optimization.
@@
There has been significant interest in using ML to accelerate the solution of integer programs. The vast majority of the work in this space focuses on learning search strategies for exact MILP solvers. This includes parameter tuning (Xu et al., 2011), preprocessing (Berthold & Hendel, 2021), branching variable selection (Khalil et al., 2016; Alvarez et al., 2017; Gasse et al., 2019; Zarpellon et al., 2021), node selection (He et al., 2014), heuristic selection (Chmiela et al., 2021), and cut selection and generation (Deza & Khalil, 2023). Another line of research in ML-for-MILP relates to learning to generate integer solutions heuristically (Nair et al., 2020; Khalil et al., 2022; Ding et al., 2020; Sonnerat et al., 2021; Song et al., 2020; Bertsimas & Stellato, 2022; Huang et al., 2023; Ye et al., ). We refer to the surveys of Bengio et al. (2021) and Zhang et al. (2023) for more details. In contrast, there has been much less work on MINLP. Illustrative examples include the work of Cauligi et al. (2021) who proposed a two-stage algorithm for quickly finding high-quality solutions for mixed-integer convex programs (MICPs), Baltean-Lugojan et al. (2019) who use supervised learning to select cuts for quadratic optimization, Nowak et al. (2018) who learn to solve quadratic assignment problems with graph networks, and Bonami et al. (2022) who use a classifier to decide on the linearization of mixed-integer quadratic problems. Most relevant to our method is the recently proposed SurCO approach of Ferber et al. (2023). They focus on mixed-integer problems with non-linear objective and linear constraints, learning to approximate the former with a linear function for a simpler heuristic optimization. Our approach differs from all of the above in its scope, addressing the most general class of MINLPs.
@@ 3 Learning to Optimize MINLPs: a Problem Formulation
$\min_{\Theta}$ $\quad\mathbb{E}\big{[}{\bf f}({\bf\hat{x}},{\bm{\xi}})\big{]},\quad\text{s.t.}\quad{\bf g}({\bf\hat{x}},{\bm{\xi}})\leq 0,\quad{\bf\hat{x}}\in\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}},\quad{\bf\hat{x}}=\bm{\psi}_{\Theta}({\bm{\xi}}).$
Here, ${\bm{\xi}}^{i}\in\mathbb{R}^{n_{\xi}}$ is a vector of instance parameters which vary across different instances; the mapping $\bm{\psi}_{\Theta}({\bm{\xi}}^{i})$ is a neural network with weights $\Theta$ that outputs a parametric solution ${\bf\hat{x}}^{i}$; ${\bf\hat{x}}^{i}=(\hat{{\bf x}}^{i}_{r},\hat{{\bf x}}^{i}_{z})$ is a predicted assignment for the mixed-integer decision variables, where $\hat{{\bf x}}^{i}_{r}\in\mathbb{R}^{n_{r}}$ and $\hat{{\bf x}}^{i}_{z}\in\mathbb{Z}^{n_{z}}$ represent the continuous and integer parts, respectively. The goal is to find the neural network weights that minimize the expected objective function ${\bf f}({\bf\hat{x}},{\bm{\xi}})$ over the parameter distribution, subject to the constraints ${\bf g}({\bf\hat{x}},{\bm{\xi}})\leq 0$. Note that $\bf{g}(\cdot)$ is a vector-valued function representing one or more inequality constraints. As is typical in MINLP, we assume that the objective and constraint functions are differentiable.
As is typical, we will train the neural network using empirical risk minimization on a sample of $m$ training instances. Then, the average value of the objective function ${\bf f}(\cdot)$ serves as a natural loss function. Our approach is self-supervised since the loss calculation does not require any labeled data. This is particularly appealing as computing optimal or even feasible solutions to a MINLP is, in general, extremely challenging. Solely minimizing the average objective is insufficient if the solutions violate the constraints. Therefore, similarly to Donti et al. (2021), we incorporate penalty terms into the loss function to account for constraint violations, enhancing the feasibility of the solution and resulting in a soft-constrained empirical risk minimization loss function given:
$\mathcal{L}(\Theta)=\frac{1}{m}\sum_{i=1}^{m}\Big{(}{\bf f}({\bf\hat{x}}^{i},{\bm{\xi}}^{i})+\lambda\cdot\left({\bf g}({\bf\hat{x}}^{i},{\bm{\xi}}^{i})\right)_{+}\Big{)}\text{ with }{\bf\hat{x}}^{i}=\bm{\psi}_{\Theta}({\bm{\xi}}^{i}),$ (2)
where $\left(\cdot\right)_{+}$ ensures only positive constraint violations are penalized (implemented via a ReLU function), and $\lambda>0$ is a penalty hyperparameter that balances the trade-off between minimizing the objective function and satisfying the constraints.
@@ 5 Learning to Optimize MINLPs with Correction Layers
Figure 1: Conceptual diagram for our self-supervised differentiable programming-based solution approach for parametric MINLP problems.
Our learnable correction layers, Rounding Classification (RC) and Learnable Threshold (LT), are designed to handle the integrality constraints of MINLPs. We decompose the mapping $\psi_{\Theta}:\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}}$ from an instance parameter vector to a candidate mixed-integer solution into two steps:
  1. 1.
  2. 2.
    The second step is a differentiable correction layer $\varphi_{\Theta_{2}}:\mathbb{R}^{n_{r}+n_{z}}\times\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}}$ that takes as input the instance parameter vector and the continuous solution produced in the first step, and outputs a candidate mixed-integer solution while maintaining differentiability. Here, $\Theta_{2}$ represents the weights of the neural network $\delta_{\Theta_{2}}:\mathbb{R}^{n_{r}+n_{z}}\times\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}+n_{z}}$, which implicitly influences the rounding strategy employed by the correction layer $\varphi_{\Theta_{2}}$.
They differ in how to determine the rounding direction but are equally easy to train with gradient descent and fast at test time. RC utilizes a classification-based stochastic rounding approach, while LT employs learnable thresholds to determine rounding directions. Further details are provided in Appendix B.
Alogorithm.
Algorithm 1 summarizes both of our approaches. Line 1 invokes the first step’s network $\pi_{\Theta_{1}}$ and lines 2–11 describe both versions of $\varphi_{\Theta_{2}}$. Our correction layers are not only simple and efficiently computable but also designed to be trainable to refine its rounding strategy. While the STE and Gumbel-Sigmoid techniques have been used to train binarized or quantized neural networks, they have not been leveraged in the context of learning-to-optimize to our knowledge. As we will see in the experimental results, the simplicity of the correction layers is key to fast solution generation in large-scale MINLP problems.
@@
1:Instance of the problem parameters ${{\bm{\xi}}^{i}}$, neural networks $\pi_{\Theta_{1}}(\cdot)$ and $\delta_{\Theta_{2}}(\cdot)$
@@
Finally, during training, the loss function eq. 2 is used to train the neural network weights $\Theta=\Theta_{1}\cup\Theta_{2}$, implicitly taking into account the objective function value and constraint violations of the predicted mixed-integer solution $\hat{\bf x}^{i}$. This process is illustrated in Figure 1. Additionally, an example of the evolution of predicted solutions during training is provided in Appendix C for further visualization.